# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 10.3 Horizontal Reservations, Adjustments and the Minimum Guarantee Choice Rule
Consider any category $v$ and two applicants, $i$ and $j$, who are both eligible for this category and satisfy the conditions in Definition 21. Awarding a position to applicant $j$ at the expense of applicant $i$ cannot be justified through meritocracy or by either of the two affirmative action policies, VR and HR. Thus, similar to the no justified envy axiom given in Definition 18 within the context of the U.S. Army’s branching process, Definition 21 also accounts for any objections applicants may raise against an allocation based on multiple assignment criteria.
Central to our application, this crucial social justice axiom fails under the SCI-AKG choice rule precisely because, with the exception of meritorious reserved candidates, applicants who declare their VR-protected categories are excluded from competing for open-category HR-protected positions. Thus, addressing the root cause of the SCI-AKG choice rule’s failures is as simple as removing this exclusion, leading directly to the 2SMG choice rule as a minimalist reform.
The simplicity of this remedy likely explains how members of the Indian judiciary were eventually able to identify the correct procedure to implement VR and HR policies concurrently, albeit after a quarter-century delay.
@@ 10.5 External Validity for Minimalist Market Design: Life Imitates Science in a Landmark Judgment
Before its publication, the initial draft of Sönmez and Yenmez (2022) was circulated in March 2019 as a Boston College working paper Sönmez and Yenmez (2019). In this work, we established that the Supreme Court-mandated SCI-AKG choice rule had been consistently undermining intended affirmative action measures in real-world practices by failing to satisfy the no justified envy axiom.116116116In Sönmez and Yenmez (2019), the SCI-AKG choice rule is referred to as the SCI-VHR choice rule, and the no justified envy axiom is referred to as elimination of justified envy. We also documented how this failure sparked numerous litigations between 1995 and 2020, causing recurring disruptions in the country’s recruitment processes.
This issue became especially problematic in jurisdictions with extensive reliance on HR policies. For example, women in states such as Andhra Pradesh, Bihar, Chhattisgarh, Madhya Pradesh, Rajasthan, Sikkim, and Uttarakhand were granted HR protections for 20–35% of positions through High Court mandates. Consequently, numerous High Court judgments in some of these states directly contradicted the Supreme Court’s ruling in Anil Kumar Gupta vs. State of U.P. (1995).
An illustrative example is the ruling of the Rajasthan High Court in Smt. Megha Shetty vs. State of Rajasthan (2013). In this case, a petitioner from the general category challenged the allocation of open-category HR-protected positions for women to women from the VR-protected category OBC. She argued that these women from OBC were not eligible for these positions unless they were selected without invoking the HR policy. Although the OBC women had higher merit scores and the State of Rajasthan had used a more compelling procedure that satisfies the no justified envy axiom, the petitioner’s case had merit under the SCI-AKG choice rule, which allows for justified envy in such situations.
@@
Motivated by this case and several others that contradicted the ruling in Anil Kumar Gupta vs. State of U.P. (1995), we proposed the 2SMG choice rule in Sönmez and Yenmez (2019) as a minimalist reform of the SCI-AKG choice rule.117117117As documented in Sönmez and Yenmez (2019, 2022), the challenges arising from the failures of the Supreme Court-mandated SCI-AKG choice rule extended far beyond conflicting with high court judgments that upheld choice rules satisfying the no justified envy axiom. In many instances, state institutions using the SCI-AKG choice rule were challenged in court by VR-protected candidates due to cases of justified envy. Some of these cases were ruled in favor of the state, while others favored the plaintiffs. Additionally, there were situations where VR-protected candidates were denied their open-category HR protections even though they did not claim their VR protections. Court decisions on these challenges also varied, sometimes favoring the candidates and other times not. Moreover, in other cases, VR-protected candidates were denied open-category HR-protected positions even when no positions were reserved for their VR-protected categories.
Following the circulation of Sönmez and Yenmez (2019), the Supreme Court revoked the SCI-AKG choice rule after 25 years in Saurav Yadav vs. State of U.P. (2020), thereby providing external validity for minimalist market design.
Over the years, two opposing views implementing the HR policy had prevailed in High Court judgments. Under the first view, exemplified by Smt. Megha Shetty vs. State of Rajasthan (2013), the High Courts of Rajasthan, Bombay, Uttarakhand, and Gujarat upheld choice rules that satisfy the no justified envy axiom. In contrast, under the second view, the High Courts of Allahabad and Madhya Pradesh enforced the SCI-AKG choice rule.
In alignment with the High Courts that adopted the first view, the justices eliminated the exclusion of VR-protected individuals from horizontal adjustments and mandated the no justified envy axiom in India through their judgment in Saurav Yadav vs. State of U.P. (2020). In paragraphs 31–33, the justices ruled:
@@
Before this judgment, under the SCI-AKG choice rule, HR-protected positions in the open category could not be awarded to members of a VR-protected category—unless they were meritorious reserved candidates—even if HR protections were not fully honored in the open category. With the removal of this restriction in Saurav Yadav vs. State of U.P. (2020), the following axiom also effectively became mandated, although it was not the primary focus of the case.
@@
Moreover, as a potential replacement for the revoked SCI-AKG choice rule, the justices in Saurav Yadav vs. State of U.P. (2020) endorsed the 2SMG choice rule as a procedure that satisfies the no justified envy axiom, even though they did not explicitly mandate it.
Central to our purposes, this reform was driven purely by the failure of a key social justice principle—the no justified envy axiom—under the SCI-AKG choice rule, a rule that had been mandated by the Supreme Court for 25 years. The justices’ primary concern was to enforce this essential principle, with their advocacy for a specific procedure being a secondary consideration. Therefore, similar to the school choice reforms for Boston, Chicago, and England (see Section 6), and the branching reform of the U.S. Army (see Section 8), the Supreme Court’s choice rule underwent reform as a direct consequence of a significant failure in the existing system based on a vital principle.
There is an additional and subtle aspect of the judgment in Saurav Yadav vs. State of U.P. (2020) that is particularly significant, even though it may not be widely understood in India. In addition to mandating the no justified envy axiom, the justices made an important clarification regarding the implementation of VR policy in the presence of HR policy.
In Indra Sawhney vs. Union of India (1992), it was established that a VR-protected position cannot be awarded to beneficiaries who qualify for an open position based on merit. However, neither this landmark judgment nor Anil Kumar Gupta vs. State of U.P. (1995)—which clarified various operational details of the HR policy—specified who qualifies for an open position when HR protections are also considered. This ambiguity was finally resolved in Saurav Yadav vs. State of U.P. (2020), where the court legally defined an individual deserving of an open position as one who qualifies for an open category position “with or without” invoking HR protections. In paragraph 36, the justices ruled:
@@
Due to this crucial clarification, the following axiom is also mandated by Saurav Yadav vs. State of U.P. (2020):
@@
    Awarding $i$ an open-category position instead, with or without replacing another individual, does not increase the number of HR-protected positions honored within the open category.118118118Prior to Saurav Yadav vs. State of U.P. (2020), unlike the first two conditions, the necessity of this third condition as a prerequisite for the assignment of VR-protected positions remained unclear.
Together with the three other axioms also mandated under Saurav Yadav vs. State of U.P. (2020)no justified envy, maximal accommodation of HR protections and non-wastefulness—this clarification has a significant consequence in India.
@@
Consider a setting in which no individual belongs to multiple HR-protected groups. Then, 2SMG is the only choice rule that satisfies non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections.
Therefore, while the justices have not explicitly mandated the 2SMG choice rule in Saurav Yadav vs. State of U.P. (2020), they have, in effect, mandated it in the country, as no other choice rule satisfies the explicit mandates of the judgment.119119119The 2SMG choice rule is mandated in the state of Gujarat under the High Court judgment Tamannaben Ashokbhai Desai vs. Shital Amrutlal Nishar (2020).
Aligned with the version of the problem reviewed in Saurav Yadav vs. State of U.P. (2020), Theorem 12 pertains to the basic version of the problem, where all positions are identical and no individual belongs to more than one HR-protected group.120120120Strictly speaking, Saurav Yadav vs. State of U.P. (2020) does not rule out the possibility of individuals belonging to multiple HR-protected groups; it simply does not address this scenario.
@@ 10.6 Overlapping Horizontal Reservations
While Persons with Disabilities (PwD) is the only HR-protected group recognized nationwide in India by a Supreme Court judgment,121121121HR protections for Persons with Disabilities were established in the Supreme Court judgment Union of India vs. National Federation of the Blind (2013). other groups, such as Women (W), receive HR protections in several states through mandates issued by High Courts or state governments. As a result, unlike VR-protected groups, individuals can often belong to multiple HR-protected groups in practical applications.
How does the possibility of overlapping HR protections influence the judgment in Saurav Yadav vs. State of U.P. (2020) and our analysis in Sections 10.3 through 10.5? We address these questions next.
Let’s start with a key observation. In India, when institutions announce positions for recruitment or admissions, they categorize these positions based on vertical categories (including open) and horizontal protections (or lack thereof). For example, in a setting with four VR-protected categories—SC, ST, OBC, and EWS—and two HR-protected groups—PwD and W—specific numbers of positions are designated for pairs such as SC-W or Open-PwD. In practice, this means that when an individual belonging to multiple HR-protected groups is awarded a position, only one of their HR protections is honored. Thus, if a woman with disabilities is awarded a position, she fulfills either the HR protection for W or for PwD, but not both.122122122This convention is referred to as one-to-one HR matching in Sönmez and Yenmez (2022). See Sönmez and Yenmez (2020) for an analysis of the alternative one-to-all HR matching convention, in which an individual fulfills all their HR protections.
This observation is significant because neither the SCI-AKG nor the minimum guarantee choice rule is well-defined under this convention in cases with overlapping HR protections. That is, without additional specifications, both procedures can yield multiple outcomes that align with their basic descriptions.
@@
If Amita receives the Open-PwD position, then the Open-W position is awarded to Chandra—the only other female applicant—in the first step. If, on the other hand, Amita receives the Open-W position, then the Open-PwD position is awarded to Debraj—the only other applicant with a disability—in the first step. In either case, Bhaskar receives the third open category position in the second step.
@@
Using this approach, the minimum guarantee choice rule can be extended with two possible orders of precedence in Example 8: either (1) the Open-PwD position is awarded first, followed by the Open-W position, with any unfilled open category positions awarded last; or (2) the Open-W position is awarded first, followed by the Open-PwD position, and any unfilled open category positions awarded last. In Example 8, positions are awarded to Amita, Chandra, and Bhaskar under the first order of precedence, and to Amita, Debraj, and Bhaskar under the second.
It could be argued that the former is better than the latter since it denies a position to the lowest merit-score individual, Debraj, rather than to the second-lowest merit-score individual, Chandra, who is denied one in the latter. Given that both allocations satisfy the minimum guarantees for both HR-protected groups, why admit a lower-merit-score individual at the expense of an individual with a higher merit score? So, in Example 8, awarding the Open-PwD position before the Open-W position results in a more “meritorious” outcome.
For the same reason, allocating the Open-W position before the Open-PwD position in this example would also result in a violation of the no justified envy axiom, as defined in Definition 21: Debraj would be awarded a position at the expense of Chandra, even though he neither has a higher merit score than she does, nor does awarding him a position at her expense increase the number of HR protections that are honored.
@@
It is worth emphasizing that, while extending the minimum guarantee choice rule via a fixed order of precedence may lead to a failure of the no justified envy axiom, it remains unclear whether this constitutes a violation of Saurav Yadav vs. State of U.P. (2020)—unlike cases involving non-overlapping HR protections.
@@
In the current setting, however, challenging an individual’s assignment may require adjustments to one or more other assignments. Specifically, although the challenge is directed at a particular individual, the correction may also involve others. In Example 8, for instance, Chandra’s potential objection to Debraj’s assignment in the first scenario would also entail a change in Amita’s assignment from Open-W to Open-PwD. While, in my opinion, such a challenge aligns with Saurav Yadav vs. State of U.P. (2020),123123123As the basis for my opinion, see Mamta Bisht vs. State of Uttaranchal (2005) at the Uttarakhand High Court, where the plaintiff successfully challenged an individual’s assignment in the open category indirectly. She argued that a third party who had been awarded an SC position should instead be assigned the open category position, thereby allowing the SC position to be reassigned to the plaintiff. as an economist, I cannot assert that this point is clear-cut under reservation jurisprudence.
As a final observation on this point, it is worth noting that for Chandra to challenge Debraj’s assignment, she would need to know Amita’s disability status. While the challenge may indeed have merit in principle under Saurav Yadav vs. State of U.P. (2020), it may rely on private information, potentially making it more difficult to pursue.
@@
There is a single category—say, the open category—with two positions. Within the open category, there are two HR-protected groups—PwD and W—each with one HR-protected position. The order of precedence is: Open-W position first, Open-PwD position second. There are three applicants—Amita, Bhaskar, and Chandra—with the following characteristics:
@@
Amita is awarded the Open-W position as the highest-merit-score woman. Since she is the only individual with a disability, the Open-PwD position is then assigned to the next highest-merit-score individual, Bhaskar, who does not have a disability. Consequently, only one of the HR-protected positions is honored in this outcome. However, if Amita were instead assigned the Open-PwD position, Chandra could be awarded the Open-W position, thereby honoring both HR-protected positions.
@@
10.6.1 Smart Reserves and Meritorious Horizontal Choice Rule

### v4

@@ 10.3 Horizontal Reservations
Consider any category $v$ and two applicants, $i$ and $j$, who are both eligible for this category and satisfy the conditions in Definition 21. Awarding a position to applicant $j$ at the expense of applicant $i$ cannot be justified either by meritocracy or by the two affirmative action policies, VR and HR (see Panel (d) of Figure 27). Thus, similar to the no justified envy axiom stated in Definition 18 within the context of the U.S. Army’s branching process, Definition 21 likewise accounts for any objections applicants may raise against an allocation based on multiple assignment criteria.
Central to our application, this crucial social justice axiom fails under the SCI–AKG choice rule precisely because, with the exception of meritorious reserved candidates, applicants who declare their VR-protected categories are excluded from competing for open-category HR-protected positions. Thus, addressing the root cause of the SCI–AKG choice rule’s failures is as simple as removing this exclusion, leading directly to the 2SMG choice rule as a minimalist reform.
For the setting in Example 12, we have already examined the 2SMG choice rule in Example 26 using its standard and computationally efficient formulation. To highlight its close connection with SCI–AKG, we now re-examine it through an alternative formulation based on horizontal adjustments.
Example 12c. As shown in Figure 28, the process is identical to SCI–AKG until the second phase of the open-category horizontal adjustment, depicted in Panel (c) (compare Panels (a) and (b) in Figures 27 and 28). At this stage, whereas SCI–AKG awards a measure of 30 adjusted positions to general-category women, 2SMG awards them to the highest-scoring unmatched women from any category. Consequently, under 2SMG, a measure of 10 general-category women and 20 category-R women with scores between 50 and 60 receive these adjusted units (contrast Panel (c) of Figure 28 with Panel (c) of Figure 27, where this “adjustment” is indicated with A in both choice rules).
After this stage, the two processes converge again, allocating the VR-protected positions using the minimum guarantee choice rule for the remaining applicants (see Panel (d) in Figures 27 and 28).
The simplicity of this minimalist remedy likely explains how members of the Indian judiciary eventually identified the correct procedure to implement VR and HR policies concurrently, albeit after a quarter-century delay.
Figure 28: Execution of the 2SMG choice rule via the horizontal adjustment process. Building on Example 21, Panels (a)–(d) trace how the 2SMG choice rule operates with horizontal adjustments to implement the minimum guarantee for the continuum economy in Example 12, illustrated in Panel (a) of Figure 26. Compared with Figure 27, the key difference is in Panel (c), where women from VR-protected categories remain eligible for positions created by the horizontal adjustment in Panel (b). Positions subject to horizontal adjustment are indicated with A in Panels (b) and (c). Abbreviations: P: Positions; W: Women; R: Vertically Reserved; $M^{G}$: General-category men; $M^{R}$ Reserved-category men; $W^{G}$: General-category women; $W^{R}$: Reserved-category women.
@@ 10.5 External Validity for Minimalist Market Design: Life Imitates Science in a Landmark Judgment
The Indian context offers another compelling case of external validity, both for the specific analysis in Sönmez and Yenmez (2019, 2022) and for minimalist market design as a paradigm, where the parallel between research and subsequent judiciary-driven reform is unusually direct. The very failure of the no justified envy axiom identified in that research—stemming from the SCI–AKG choice rule’s denial of open-category HR protections for VR-protected candidates—materialized in practice, causing persistent disruption for years. After decades of litigation, the Supreme Court in Saurav Yadav (2020) adopted the same minimalist correction proposed in the research, targeting the root cause by removing this denial, and replaced the SCI–AKG choice rule after 25 years.
Before its publication, the initial draft of Sönmez and Yenmez (2022) was circulated in March 2019 as a Boston College working paper Sönmez and Yenmez (2019). In this work, we established that the Supreme Court-mandated SCI–AKG choice rule had been consistently undermining intended affirmative action measures in real-world practices by failing to satisfy the no justified envy axiom.130130130In Sönmez and Yenmez (2019), the SCI–AKG choice rule is referred to as the SCI–VHR choice rule, and the no justified envy axiom is referred to as elimination of justified envy. We also documented how this failure sparked numerous litigations between 1995 and 2020, causing recurring disruptions in the country’s recruitment processes.
This issue became especially problematic in jurisdictions with extensive reliance on HR policies. For example, women in states such as Andhra Pradesh, Bihar, Chhattisgarh, Madhya Pradesh, Rajasthan, Sikkim, and Uttarakhand were granted HR protections for 20–35% of positions through High Court mandates. Consequently, numerous High Court judgments in some of these states directly contradicted the Supreme Court’s ruling in Anil Kumar Gupta (1995).
An illustrative example is the ruling of the Rajasthan High Court in Smt. Megha Shetty vs. State of Rajasthan (2013). In this case, a petitioner from the general category challenged the allocation of open-category HR-protected positions for women to women from the VR-protected category OBC. She argued that these women from OBC were not eligible for these positions unless they were selected without invoking the HR policy. Although the OBC women had higher merit scores and the State of Rajasthan had used a more compelling procedure that satisfies the no justified envy axiom, the petitioner’s case had merit under the SCI–AKG choice rule, which allows for justified envy in such situations.
@@
Motivated by this case and several others that contradicted the ruling in Anil Kumar Gupta (1995), we proposed the 2SMG choice rule in Sönmez and Yenmez (2019) as a minimalist reform of the SCI–AKG choice rule.131131131As documented in Sönmez and Yenmez (2019, 2022), the challenges arising from the failures of the Supreme Court-mandated SCI–AKG choice rule extended far beyond conflicting with high court judgments that upheld choice rules satisfying the no justified envy axiom. In many instances, state institutions using the SCI–AKG choice rule were challenged in court by VR-protected candidates due to cases of justified envy. Some of these cases were ruled in favor of the state, while others favored the plaintiffs. Additionally, there were situations where VR-protected candidates were denied their open-category HR protections even though they did not claim their VR protections. Court decisions on these challenges also varied, sometimes favoring the candidates and other times not. Moreover, in other cases, VR-protected candidates were denied open-category HR-protected positions even when no positions were reserved for their VR-protected categories.
Following the circulation of Sönmez and Yenmez (2019), the Supreme Court revoked the SCI–AKG choice rule after 25 years in Saurav Yadav (2020), thereby providing external validity for minimalist market design.
Over the years, two opposing views implementing the HR policy had prevailed in High Court judgments. Under the first view, exemplified by Smt. Megha Shetty vs. State of Rajasthan (2013), the High Courts of Rajasthan, Bombay, Uttarakhand, and Gujarat upheld choice rules that satisfy the no justified envy axiom. In contrast, under the second view, the High Courts of Allahabad and Madhya Pradesh enforced the SCI–AKG choice rule.
In alignment with the High Courts that adopted the first view, the justices eliminated the exclusion of VR-protected individuals from horizontal adjustments and mandated the no justified envy axiom in India through their judgment in Saurav Yadav (2020). In paragraphs 31–33, the justices ruled:
@@
Before this judgment, under the SCI–AKG choice rule, HR-protected positions in the open category could not be awarded to members of a VR-protected category—unless they were meritorious reserved candidates—even if HR protections were not fully honored in the open category. With the removal of this restriction in Saurav Yadav (2020), the following axiom also effectively became mandated, although it was not the primary focus of the case.
@@
Moreover, as a potential replacement for the revoked SCI–AKG choice rule, the justices in Saurav Yadav (2020) endorsed the 2SMG choice rule as a procedure that satisfies the no justified envy axiom, even though they did not explicitly mandate it.
Central to our purposes, this reform was driven purely by the failure of a key social justice principle—the no justified envy axiom—under the SCI–AKG choice rule, a rule that had been mandated by the Supreme Court for 25 years. The justices’ primary concern was to enforce this essential principle, with their advocacy for a specific procedure being a secondary consideration. Therefore, similar to the school choice reforms for Boston, Chicago, and England (see Section 6), and the branching reform of the U.S. Army (see Section 8), the Supreme Court’s choice rule underwent reform as a direct consequence of a significant failure in the existing system based on a vital principle.
There is an additional and subtle aspect of the judgment in Saurav Yadav (2020) that is particularly significant, even though it may not be widely understood in India. In addition to mandating the no justified envy axiom, the justices made an important clarification regarding the implementation of VR policy in the presence of HR policy.
In Indra Sawhney (1992), it was established that a VR-protected position cannot be awarded to beneficiaries who qualify for an open position based on merit. However, neither this landmark judgment nor Anil Kumar Gupta (1995)—which clarified various operational details of the HR policy—specified who qualifies for an open position when HR protections are also considered. This ambiguity was finally resolved in Saurav Yadav (2020), where the court legally defined an individual deserving of an open position as one who qualifies for an open category position “with or without” invoking HR protections. In paragraph 36, the justices ruled:
@@
Due to this crucial clarification, the following axiom is also effectively mandated by Saurav Yadav (2020):
@@
    Awarding $i$ an open-category position instead, with or without replacing another individual, does not increase the number of HR-protected positions honored within the open category.132132132Prior to Saurav Yadav (2020), unlike the first two conditions, the necessity of this third condition as a prerequisite for the assignment of VR-protected positions remained unclear.
Together with the three other axioms also mandated under Saurav Yadav (2020)—no justified envy, maximal accommodation of HR protections and non-wastefulness—this clarification has a significant consequence in India.
@@
Consider a setting in which no individual belongs to multiple HR-protected groups. Then, 2SMG is the unique choice rule that satisfies non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections.
Therefore, while the justices have not explicitly mandated the 2SMG choice rule in Saurav Yadav (2020), they have, in effect, mandated it in the country, as no other choice rule satisfies the explicit mandates of the judgment.133133133The 2SMG choice rule is mandated in the state of Gujarat under the High Court judgment Tamannaben Ashokbhai Desai vs. Shital Amrutlal Nishar (2020).
Aligned with the version of the problem reviewed in Saurav Yadav (2020), Theorem 12 pertains to the basic version of the problem, where all positions are identical and no individual belongs to more than one HR-protected group.134134134Strictly speaking, Saurav Yadav (2020) does not rule out the possibility of individuals belonging to multiple HR-protected groups; it simply does not address this scenario.
@@ 10.6 Overlapping Horizontal Reservations
While Persons with Disabilities (PwD) is the only HR-protected group recognized nationwide in India by a Supreme Court judgment,135135135HR protections for Persons with Disabilities were established in the Supreme Court judgment Union of India vs. National Federation of the Blind (2013). other groups, such as Women (W), receive HR protections in several states through mandates issued by High Courts or state governments. As a result, unlike VR-protected groups, individuals can often belong to multiple HR-protected groups in practical applications.
How does the possibility of overlapping HR protections influence the judgment in Saurav Yadav (2020) and our analysis in Sections 10.3 through 10.5? We address these questions next.
Let’s start with a key observation. In India, when institutions announce positions for recruitment or admissions, they categorize these positions based on vertical categories (including open) and horizontal protections (or lack thereof). For example, in a setting with four VR-protected categories—SC, ST, OBC, and EWS—and two HR-protected groups—PwD and W—specific numbers of positions are designated for pairs such as SC-W or Open–PwD. In practice, this means that when an individual belonging to multiple HR-protected groups is awarded a position, only one of their HR protections is honored. Thus, if a woman with disabilities is awarded a position, she fulfills either the HR protection for W or for PwD, but not both.136136136This convention is referred to as one-to-one HR matching in Sönmez and Yenmez (2022). See Sönmez and Yenmez (2020) for an analysis of the alternative one-to-all HR matching convention, in which an individual fulfills all their HR protections.
This observation is significant because neither the SCI–AKG nor the minimum guarantee choice rule is well-defined under this convention in cases with overlapping HR protections. That is, without additional specifications, both procedures can yield multiple outcomes that align with their basic descriptions.
@@
If Amita receives the Open–PwD position, then the Open–W position is awarded to Chandra—the only other female applicant—in the first step. If, on the other hand, Amita receives the Open–W position, then the Open–PwD position is awarded to Debraj—the only other applicant with a disability—in the first step. In either case, Bhaskar receives the third open category position in the second step (see Figure 29).
Figure 29: Allocation of HR-protected positions under overlapping horizontal reservations. For the setting in Example 13, Panel (a) shows eligibility in the first phase of the minimum guarantee choice rule for the Open category (indicated by light unbroken lines), which requires membership in an HR-protected group, and in the second phase (indicated by light broken lines), which does not. In Panels (b) and (c), assignments are indicated by bold versions of these lines. Because they each belong to at least one HR-protected group, Amita, Chandra, and Debraj are eligible in the first phase. As the highest-scoring applicant, Phase 1 starts with Amita; since she belongs to both HR-protected groups W and PwD, she may be assigned to either of these positions. If Amita is assigned the PwD position, then the W position is assigned to the only other woman, Chandra, in the first phase, and the remaining open position is assigned to the highest-scoring remaining applicant, Bhaskar, in the second phase (Panel (b)). If instead Amita is assigned the W position, then the PwD position is assigned to the only other person with a disability, Debraj, in the first phase, and again the remaining open position is assigned to Bhaskar in the second phase (Panel (c)). Abbreviations: W: Women; M: Men; PwD: Persons with Disabilities.
@@
Using this approach, the minimum guarantee choice rule can be extended with two possible orders of precedence in Example 13: either (1) the Open–PwD position is awarded first, followed by the Open–W position, with any unfilled open category positions awarded last; or (2) the Open–W position is awarded first, followed by the Open–PwD position, and any unfilled open category positions awarded last. In Example 13, positions are allocated to Amita, Chandra, and Bhaskar under the first scenario, and to Amita, Debraj, and Bhaskar under the second, as depicted in Panels (b) and (c) of Figure 29, respectively.
It could be argued that the first scenario is better than the latter since it denies a position to the lowest merit-score individual, Debraj, rather than to the second-lowest merit-score individual, Chandra, who is denied one in the latter. Given that both allocations satisfy the minimum guarantees for both HR-protected groups, why admit a lower-merit-score individual at the expense of an individual with a higher merit score? So, in Example 13, awarding the Open–PwD position before the Open–W position results in a more “meritorious” outcome.
For the same reason, allocating the Open–W position before the Open–PwD position in this example would also result in a violation of the no justified envy axiom, as defined in Definition 21: Debraj would be awarded a position at the expense of Chandra, even though he neither has a higher merit score than she does, nor does awarding him a position at her expense increase the number of HR protections that are honored.
@@
It is worth emphasizing that, while extending the minimum guarantee choice rule via a fixed order of precedence may lead to a failure of the no justified envy axiom, it remains unclear whether this constitutes a violation of Saurav Yadav (2020)—unlike cases involving non-overlapping HR protections.
@@
In the current setting, however, challenging an individual’s assignment may require adjustments to one or more other assignments. Specifically, although the challenge is directed at a particular individual, the correction may also involve others. In Example 13, for instance, Chandra’s potential objection to Debraj’s assignment in the first scenario would also entail a change in Amita’s assignment from Open–W to Open–PwD (see Figure 29). While, in my opinion, such a challenge aligns with Saurav Yadav (2020),137137137As the basis for my opinion, see Mamta Bisht vs. State of Uttaranchal (2005) at the Uttarakhand High Court, where the plaintiff successfully challenged an individual’s assignment in the open category indirectly. She argued that a third party who had been awarded an SC position should instead be assigned the open category position, thereby allowing the SC position to be reassigned to the plaintiff. as an economist, I cannot assert that this point is clear-cut under reservation jurisprudence.
As a final observation on this point, it is worth noting that for Chandra to challenge Debraj’s assignment, she would need to know Amita’s disability status. While the challenge may indeed have merit in principle under Saurav Yadav (2020), it may rely on private information, potentially making it more difficult to pursue.
@@
There is a single category—say, the open category—with two positions. Within the open category, there are two HR-protected groups—PwD and W—each with one HR-protected position. The order of precedence is: Open–W position first, Open–PwD position second. There are three applicants—Amita, Bhaskar, and Chandra—with the following characteristics:
@@
Amita is awarded the Open–W position as the highest-merit-score woman. Since she is the only individual with a disability, the Open–PwD position is then assigned to the next highest-merit-score individual, Bhaskar, who does not have a disability. Consequently, only one of the HR-protected positions is honored in this outcome. However, if Amita were instead assigned the Open–PwD position, Chandra could be awarded the Open–W position, thereby honoring both HR-protected positions (see Figure 30).
Figure 30: Allocation of HR-protected positions under overlapping horizontal reservations with a fixed order of precedence. For the setting in Example 14, Panel (a) shows eligibility in the first phase of the minimum guarantee choice rule (indicated by light unbroken lines), which requires membership in an HR-protected group, and in the second phase (indicated by light broken lines), which does not. In Panels (b) and (c), assignments are indicated by bold versions of these lines. Panel (b) presents the order of precedence, where the Open–W position is processed before the Open–PwD position: Amita receives the Open–W position and Bhaskar the Open–PwD position, so only the HR-protection for W is honored. Panel (c) presents the counterfactual order of precedence, where the Open–PwD position is processed first: Amita is assigned the Open–PwD position and Chandra the Open–W position, thereby honoring both HR protections. Abbreviations: W: Women; M: Men; PwD: Persons with Disabilities.
@@
10.6.1 Smart Reserves and the Two-Step Meritorious Horizontal Choice Rule
We introduce a generalization of the minimum-guarantee choice rule through a smart reserve system and utilize it to design a choice rule that extends 2SMG to settings with overlapping HR protections.